Ask
your own question, for FREE!
Probability
12 Online
P(dry followed by dry)=0.85 P(dry followed by wet)=0.15 Probability of dry day is 0.80 Question: whats is the Probability(wet day followed by another wet day)??
Still Need Help?
Join the QuestionCove community and study together with friends!
i really need help with this assignment question
prob of wet day 1-0.8 = 0.2 probability of wet day followed by wet day 0.2*0.2=0.04
If P(dry)=0.8, then P(dry=>dry) should be 0.8^2=0.64 ...
im looking at both yours answers and working one says 0.04 and the other 0.64 hmmmm
No I'm not the answer, I'm just doubting the validity of the question.
Still Need Help?
Join the QuestionCove community and study together with friends!
ye the question just doesnt give out much info which isnt any help
what ya mean by validity?? @kc_kennylau
The first statement and the third statement contradict each other @divu.mkr
yep...
good point :) @kc_kennylau
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
3 hours ago
0 Replies
0 Medals
3 days ago
3 Replies
2 Medals
6 days ago
3 Replies
0 Medals
1 day ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal